OpenStudy (anonymous):

A bowl contains candies of the same size in three flavors: orange, strawberry, and pineapple. If the probability of randomly pulling out an orange candy is fraction 2 over 8, and the probability of randomly pulling out a strawberry candy is fraction 3 over 8, what is the probability of randomly pulling out a pineapple candy?

OpenStudy (anonymous):

2/8+3/8 =5/8?

OpenStudy (anonymous):

2/8-3/8 = 1/8?

OpenStudy (anonymous):

64% is what i got and ill show you how...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We need to end up with \(\Large\rm \frac{8}{8}\) with our three individual odds.

OpenStudy (anonymous):

So far we have \(\Large\rm \frac{2}{8}\) and \(\Large\rm \frac{3}{8}\)

OpenStudy (anonymous):

o then the pineapple would be 3/8

OpenStudy (anonymous):

since they should be equal?

OpenStudy (anonymous):

so im sorry i meant 3/8 not 64%-i dont know where that came from :b

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

thanks @Kim21 got it

OpenStudy (anonymous):

welcome :)